# On the Hardness of Approximating Distributions with Probabilistic Circuits

### v1

@@ Appendix B Relationship between Approximate Modeling and Inference
We have shown that even if we allow some approximation error, it is hard to efficiently approximate distributions using tractable probabilistic circuits. As approximating modeling is a hard task, one would hope that ensuring approximate modeling would guarantee the approximation of hard inference queries with bounded error. Unfortunately, this is not the case for all notions of approximation. Therefore, we study the relationship between approximate modeling and inference, in particular focusing on absolute approximation of marginal, conditional, and maximum-a-posteriori (MAP) queries.
@@
In Section 3, we showed that bounded total variation distance is a necessary but insufficient condition for relative approximation of marginals. We now consider a slightly weaker notion of approximation called absolute approximation. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. Then we say another distribution $Q(\mathbf{X})$ is an absolute approximator of marginals of $P$ w.r.t. $0\leq\epsilon\leq 1$ if: $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert\leq\epsilon$ for every assignment $\mathbf{y}$ to a subset of variables $Y\subseteq\mathbf{X}$. We show that any model that is a $k\epsilon^{2}$-$D_{f}(P\|Q)$-approximator of $P$ must also be an absolute approximator of marginals of $P$ w.r.t. $\epsilon$.
Theorem B.1 (Bounded $D_{f}$ implies absolute approx. of marginals).
@@
Given two distributions $P(\mathbf{X})$ and $Q(\mathbf{X})$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{f}(P\|Q)<k\epsilon^{2}$ then for all assignments $\mathbf{y}$ to a subset $\mathbf{Y}\subseteq\mathbf{X}$, $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$.
@@
Note that while the absolute error of marginals is symmetric between $P$ and $Q$, $f$-divergence between $P$ and $Q$, such as the KL-divergence, is not symmetric. Therefore, we utilize the implications derived in [27], that $D_{f}(P\|Q)<k\epsilon^{2}$ then $D_{\mathsf{TV}}(P\|Q)<\epsilon$. Moreover, given that the total variation distance is an $f$-divergence, we know that by the monotonicity property $D_{f}(P(\mathbf{Y}),Q(\mathbf{Y}))\leq D_{f}(P(\mathbf{X}),Q(\mathbf{X}))$ [29] as $\mathbf{Y}\subseteq\mathbf{X}$. By definition, $\max_{S\subseteq\{0,1\}^{n}}\left\lvert P(S)-Q(S)\right\rvert<\epsilon$. Since these are discrete probability distributions, the event $S=\mathbf{y}$ is valid. Therefore, $\forall\mathbf{y}:\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$. ∎
From the above proof, we immediately derive the following corollary.
Corollary B.2 (Bounded $D_{\mathsf{TV}}$ implies absolute approx. of marginals).
Given two distributions $P(\mathbf{X})$ and $Q(\mathbf{X})$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{\mathsf{TV}}(P\|Q)<\epsilon$, then for all assignments $\mathbf{y}$ to $\mathbf{Y}\subseteq\mathbf{X}$, $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$.
Since marginals are tractable for decomposable PCs, approximating a target distribution with bounded $f$-divergence using a decomposable PC implies that marginals can be approximated in polynomial-time with bounded absolute error. Previously, Dagum and Luby [12] considered this query for Bayesian networks and showed that there exists a randomized polynomial-time algorithm for the absolute approximation of marginals. Due to this, it is unsurprising that this form of approximation is guaranteed by the more restrictive condition of bounded f-divergence.
Moving to more difficult queries, we next study approximate inference implications for MAP. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. We say another distribution $Q(\mathbf{X})$ is an absolute approximator of the maximum-a-posteriori of $P$ w.r.t. $0\leq\epsilon\leq 1$ if: for every assignment $\mathbf{e}$ (called the evidence) to a subset $\mathbf{E}\subseteq\mathbf{X}$, $\left\lvert\max_{\mathbf{y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}}Q(\mathbf{y},\mathbf{e})\right\rvert\leq\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$. We next show that $k\epsilon^{2}$-$D_{{f}}$-approximators are also absolute approximators of MAP w.r.t. $\epsilon$.
Theorem B.3 (Bounded $D_{f}$ implies absolute approx. of MAP).
Given two distributions $P(\mathbf{X})$ and $Q(\mathbf{X})$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{f}(P\|Q)<k\epsilon^{2}$ then for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$, $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$.
@@
Analogous to Theorem B.1, we utilize the fact that if we have $D_{f}(P\|Q)<k\epsilon^{2}$, we know $D_{\mathsf{TV}}(P\|Q)<\epsilon$. Using $D_{\mathsf{TV}}(P\|Q)<\epsilon$, we have $\max_{\mathbf{x}}\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert<\epsilon$ by definition. Then for all $\mathbf{x}$, $Q(\mathbf{x})-\epsilon<P(\mathbf{x})<Q(\mathbf{x})+\epsilon$. W.l.o.g., suppose $\max P(\mathbf{x})>\max Q(\mathbf{x})$. Then $\left\lvert\max P(\mathbf{x})-\max Q(\mathbf{x})\right\rvert<(\max Q(\mathbf{x})+\epsilon)-\max Q(\mathbf{x})=\epsilon$. Thus, as this was for all $\mathbf{X}$, we can extend this to subsets of $\mathbf{X}$. Finding that for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$, $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$. Thus, approximating with bounded $f$-divergence by a deterministic PC implies polynomial-time approximation of MAP with bounded absolute error. ∎
Again, we restrict this to the special case of total variation distance.
Corollary B.4 (Bounded $D_{\mathsf{TV}}$ implies absolute approx. of MAP).
Given two distributions $P(\mathbf{X})$ and $Q(\mathbf{X})$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{\mathsf{TV}}(P\|Q)<\epsilon$ then for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$, $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$.
Thus, a deterministic PC that is an $\epsilon$-$D_{\mathsf{TV}}$-approximator of a distribution $P$ would imply that exact MAP inference w.r.t. this PC grants us tractable approximate MAP inference w.r.t. the original model. However, the converse does not hold: a PC that can be used for approximate MAP inference is not necessarily a good approximation of the full distribution. In particular, we can obtain PCs with identical MAP solutions, but result in large total variation distance due to disjoint support, as seen in Counterexample 1.
Lastly, not all tractable queries for PCs are guaranteed to admit absolute approximation even under the restrictive framework of approximate modeling with bounded distance.
Theorem B.5 (Bounded $D_{\mathsf{TV}}$ does not imply absolute approx. of conditionals/conditional MAP).
@@
  • Proof sketch.
    Let $P$ be a probability distribution over $\mathbf{X}$, and let some assignment $\mathbf{e}$ to $\mathbf{E}\subseteq\mathbf{X}$ be evidence. We construct a distribution $Q$ using the mass function $Q(\mathbf{y}^{*},\mathbf{e})=P(\mathbf{e})P(\mathbf{y}^{*}|\mathbf{e})+k\epsilon P(\mathbf{e})$ where $\mathbf{y}^{*}$ maximizes $P(\mathbf{y}|\mathbf{e})$. Then we remove the mass we added to ensure $Q$ is still a valid probability distribution. This allows us to construct a distribution with $D_{\mathsf{TV}}(P\|Q)<\epsilon$ but conditional MAP equals $k\epsilon$ where $k$ can be arbitrarily large. We refer to Appendix E.6 for the full distribution construction and detailed proof. ∎
Moreover, because relative approximation of conditionals imply their absolute approximation [12], bounded $D_{\mathsf{TV}}$ also does not imply relative approximation of conditionals. However, it is well known that both absolute and relative approximation of conditionals is $\mathsf{NP}$-hard in Bayesian networks [12]. Even though approximate modeling with a bounded $D_{\mathsf{TV}}$ is also an $\mathsf{NP}$-hard task, it still does not guarantee a polynomial-time algorithm for approximating conditional queries. This highlights a key limitation: while tractability of queries is guaranteed by the structural properties of our learned PCs, some queries do not yield “good” approximations for all assignments even after learning within bounded distance.
Appendix C $k$-convex $f$-divergences
We rewrite the table from [27] here.
@@ Appendix C $k$-convex $f$-divergences
relative entropy (KL) $t\log t$ $\frac{1}{M}$ $(0,M]$
total variation $\frac{|t-1|}{2}$ 00 $(0,\infty)$
Pearson’s $\chi^{2}$ $(t-1)^{2}$ $2$ $(0,\infty)$
@@
reverse relative entropy $-\log t$ $\frac{1}{M^{2}}$ $(0,M]$
@@
Table 1: Examples of Strongly Convex Divergences.
Appendix D Rectangle Partitions

### v2

@@ 5 Relationship between Approximate Modeling and Inference
We have shown that even if we allow some approximation error, it is hard to efficiently approximate distributions using tractable probabilistic circuits. Given that approximate modeling remains a hard task, one would hope that computing the approximators with bounded distance would allow us to approximate hard inference queries with bounded error. In this section, we study the relationship between approximate modeling and inference, in particular focusing on relative and absolute approximations of marginal, conditional, and maximum-a-posteriori (MAP) queries.
@@
In Section 3.1, we showed that bounded total variation distance is a necessary but not sufficient condition for relative approximation of marginals. We now consider a slightly weaker notion of approximation called absolute approximation. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. Then we say another distribution $Q(\mathbf{X})$ is an absolute approximator of marginals of $P$ with respect to $0\leq\epsilon\leq 1$ if: $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert\leq\epsilon$ for every assignment $\mathbf{y}$ to a subset of variables $Y\subseteq\mathbf{X}$. We show that any model that is a $k\epsilon^{2}$-$D_{f}$-approximator of $P$ must also be an absolute approximator of marginals of $P$ with respect to $\epsilon$.
Theorem 5.1 (Bounded $D_{f}$ implies absolute approximation of marginals).
@@
Given two distributions $P$ and $Q$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{f}(P\|Q)<k\epsilon^{2}$ then for all assignments $\mathbf{y}$ to a subset $\mathbf{Y}\subseteq\mathbf{X}$, we have $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$.
@@
Note that while the absolute error of marginals is symmetric between $P$ and $Q$, $f$-divergence between $P$ and $Q$, such as the KL-divergence, is not symmetric. Therefore, we utilize the implications derived in [34], that $D_{f}(P\|Q)<k\epsilon^{2}$ then $D_{\mathsf{TV}}(P\|Q)<\epsilon$. Moreover, given that the total variation distance is an $f$-divergence, we know that by the monotonicity property [36] $D_{f}(P(\mathbf{Y}),Q(\mathbf{Y}))\leq D_{f}(P(\mathbf{X}),Q(\mathbf{X}))$ for any $\mathbf{Y}\subseteq\mathbf{X}$. By definition, $\max_{S\subseteq\{0,1\}^{n}}\left\lvert P(S)-Q(S)\right\rvert<\epsilon$, and thus $\forall\mathbf{y}:\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$. ∎
From the above proof, we also immediately derive the following corollary.
Corollary 5.2 (Bounded $D_{\mathsf{TV}}$ implies absolute approximation of marginals).
Given two distributions $P$ and $Q$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{\mathsf{TV}}(P\|Q)<\epsilon$, then for all assignments $\mathbf{y}$ to $\mathbf{Y}\subseteq\mathbf{X}$, we have $\left\lvert P(\mathbf{y})-Q(\mathbf{y})\right\rvert<\epsilon$.
Since marginals are tractable for decomposable PCs, approximating a target distribution with bounded $f$-divergence using a decomposable PC implies that marginals can be approximated in polynomial-time with bounded absolute error. This aligns with Dagum and Luby [14], who showed that there exists a randomized polynomial-time algorithm for the absolute approximation of marginals for Bayesian networks.
We next study approximate inference implications for MAP inference. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. We say another distribution $Q(\mathbf{X})$ is an absolute approximator of the maximum-a-posteriori of $P$ with respect to $0\leq\epsilon\leq 1$ if: for every assignment $\mathbf{e}$ (called the evidence) to a subset $\mathbf{E}\subseteq\mathbf{X}$, $\left\lvert\max_{\mathbf{y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}}Q(\mathbf{y},\mathbf{e})\right\rvert\leq\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$. We next show that $k\epsilon^{2}$-$D_{{f}}$-approximators are also absolute approximators of MAP with respect to $\epsilon$.
Theorem 5.3 (Bounded $D_{f}$ implies absolute approximation of MAP).
Given two distributions $P$ and $Q$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{f}(P\|Q)<k\epsilon^{2}$ then for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$, we have $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$.
@@
Analogous to Theorem 5.1, we utilize the fact that if we have $D_{f}(P\|Q)<k\epsilon^{2}$, we know $D_{\mathsf{TV}}(P\|Q)<\epsilon$. Using $D_{\mathsf{TV}}(P\|Q)<\epsilon$, we have $\max_{\mathbf{x}}\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert<\epsilon$ by definition. Then for all $\mathbf{x}$, $Q(\mathbf{x})-\epsilon<P(\mathbf{x})<Q(\mathbf{x})+\epsilon$. W.l.o.g., suppose $\max P(\mathbf{x})>\max Q(\mathbf{x})$. Then $\left\lvert\max P(\mathbf{x})-\max Q(\mathbf{x})\right\rvert<(\max Q(\mathbf{x})+\epsilon)-\max Q(\mathbf{x})=\epsilon$. Thus, as this holds for all $\mathbf{x}$, we can extend this to $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$ and $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$. Thus, approximating with bounded $f$-divergence by a deterministic and decomposable PC implies polynomial-time approximation of MAP with bounded absolute error. ∎
Again, we can restrict this to the special case of total variation distance via the above.
Corollary 5.4 (Bounded $D_{\mathsf{TV}}$ implies absolute approximation of MAP).
Given two distributions $P$ and $Q$ over a set of variables $\mathbf{X}$ and $0\leq\epsilon\leq 1$, if $D_{\mathsf{TV}}(P\|Q)<\epsilon$ then for every assignment $\mathbf{e}$ to a subset $\mathbf{E}\subseteq\mathbf{X}$, we have $\lvert\max_{\mathbf{y}\in\mathbf{Y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}\in\mathbf{Y}}Q(\mathbf{y},\mathbf{e})\rvert<\epsilon$ where $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$.
Thus, a deterministic and decomposable PC that is an $\epsilon$-$D_{\mathsf{TV}}$-approximator of a distribution $P$ would imply that exact MAP inference w.r.t. this PC grants us tractable approximate MAP inference w.r.t. the original distribution $P$. However, the converse does not hold: a PC that can be used for approximate MAP inference is not necessarily a good approximation of the full distribution.
Counterexample 1.
Consider a family of distributions $P(\mathbf{x})$ such that $\max_{\mathbf{x}}P(\mathbf{x})<\epsilon$. Then, we construct a distribution $P^{\prime}(\mathbf{X},Z)$ such that $P(\mathbf{x},Z=1)=P(\mathbf{x})$ and $P(\mathbf{x},Z=0)=0$. Similarly, let $Q(\mathbf{X},Z)$ be such that $Q(\mathbf{x},Z=0)=P(\mathbf{x})$ and $Q(\mathbf{x},Z=1)=0$. Thus, for any assignment $\mathbf{e}$ to $\mathbf{E}\subseteq\mathbf{X}\cup\{Z\}$, $\left\lvert\max_{\mathbf{y}}P(\mathbf{y},\mathbf{e})-\max_{\mathbf{y}}Q(\mathbf{y},\mathbf{e})\right\rvert\leq P(\mathbf{x})<\epsilon$, so $Q$ is an absolute approximator of the MAP of $P$. However, $D_{\mathsf{TV}}(P\|Q)=1$ as $P$ and $Q$ have disjoint supports.
Lastly, not all tractable queries for PCs are guaranteed to admit absolute approximation even under this framework of approximate modeling with bounded distance.
Theorem 5.5 (Bounded $D_{\mathsf{TV}}$ does not imply absolute approx. of conditionals/conditional MAP).
@@
Proof.
Let $P$ be a probability distribution over $\mathbf{X}$ such that $P(\mathbf{e})<1/k$ for some assignment $\mathbf{e}$ to $\mathbf{E}\subseteq\mathbf{X}$. Let $\mathbf{Y}=\mathbf{X}\setminus\mathbf{E}$. We construct another distribution $Q$ such that: $Q(\mathbf{y}^{*},\mathbf{e})=P(\mathbf{e})P(\mathbf{y}^{*}|\mathbf{e})+k\epsilon P(\mathbf{e})$ where $\mathbf{y}^{*}$ maximizes $P(\mathbf{y}|\mathbf{e})$; $Q(\mathbf{y}_{1},\mathbf{e})=P(\mathbf{e})P(\mathbf{y}_{1}|\mathbf{e})-k\epsilon P(\mathbf{e})$ for another assignment $\mathbf{y}_{1}$; and $Q(\mathbf{x})=P(\mathbf{x})$ for all other assignments $\mathbf{x}$. Note that $Q$ is normalized by construction. Then the total variation distance between $P$ and $Q$ is: $D_{\mathsf{TV}}(P\|Q)=\frac{1}{2}(\left\lvert P(\mathbf{y}^{*},\mathbf{e})-Q(\mathbf{y}^{*},\mathbf{e})\right\rvert+\left\lvert P(\mathbf{y}_{1},\mathbf{e})-Q(\mathbf{y}_{1},\mathbf{e})\right\rvert)=k\epsilon P(\mathbf{e})<\epsilon$. On the other hand, the absolute approximation error of conditional MAP can grow arbitrarily by increasing $k$: $\left\lvert\max_{\mathbf{y}}P(\mathbf{y}|\mathbf{e})-\max_{\mathbf{y}}Q(\mathbf{y}|\mathbf{e})\right\rvert=\left\lvert P(\mathbf{y}^{*}|\mathbf{e})-P(\mathbf{y}^{*}|\mathbf{e})-k\epsilon\right\rvert=k\epsilon$. Note that the same also holds for absolute approximation of conditionals. ∎
Moreover, because relative approximation of conditionals imply their absolute approximation [14], bounded $D_{\mathsf{TV}}$ also does not imply relative approximation of conditionals. It is well known that absolute and relative approximation of conditionals is $\mathsf{NP}$-hard in Bayesian networks [14]. Even though approximate modeling with a bounded $D_{\mathsf{TV}}$ is also an $\mathsf{NP}$-hard task, solving it still does not guarantee a polynomial-time algorithm for approximating conditional queries. This highlights a key limitation: while tractability of queries is guaranteed by the structural properties of our learned PCs, some queries do not yield “good” approximations for all assignments even after learning within bounded distance.
6 Conclusion and Discussions
We established the hardness of approximating distributions with tractable probabilistic models such that the $f$-divergence is small. First, we showed that this task is $\mathsf{NP}$-hard for any model supporting tractable marginal inference, including decomposable PCs. Then, we used the Sauerhoff function to demonstrate an exponential size gap between the class of decomposable PCs and that of deterministic and decomposable PCs when allowing for a bounded approximation error. This proves that the expressive efficiency gap that exists in exact compilation persists even under the relaxed approximation conditions. Finally, we characterized which queries remain well-approximated under the framework of approximate compilation.
These results highlight key challenges in learning compact and expressive PCs while maintaining tractable inference. In light of this, we ask: can a polynomial-time algorithm enable learning an $\epsilon$-approximator for a broad family of distributions with a more relaxed $\epsilon$? While this is trivial when total variation is near 1, investigating whether structurally constrained PCs remain expressive under weaker approximation could reveal key limits of learnability. Furthermore, does there exist modeling conditions that are sufficient to guarantee relative approximation of various queries? Lastly, we see this work as a first step to encourage further theoretical studies on approximate modeling and inference with guarantees using tractable models. In this paper, we focused on PCs that are tractable for marginal and MAP queries, but there are large classes of tractable models whose efficient approximation power remain largely unknown.
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their feedback towards improving this paper. We also thank Adrian Ciotinga for helpful feedback and discussion. This work was supported in part by the AFOSR grant FA9550-25-1-0320.
Appendix A Appendix
A.1 $k$-convex $f$-divergences
Table 1: Examples of $k$-convex $f$-divergences [34]
@@ A.1 $k$-convex $f$-divergences
Kullback-Leibler $t\log t$ $\frac{1}{M}$ $(0,M]$
Total variation $\frac{|t-1|}{2}$ 0 $(0,\infty)$
Pearson’s $\chi^{2}$ $(t-1)^{2}$ 22 $(0,\infty)$
@@
Reverse KL $\log t$ $\frac{1}{M^{2}}$ $(0,M]$
@@
A.2 Rectangle Partitions